OpenStudy (gigi1103):

Question: A football team lost 9 yards on its first play. On the next two plays they gained 3 yards and then 6 yards. A. What expression can be written to represent how many yards the football team is from where it started? Explain your reasoning. B. How many yards is the football team from where it started?

OpenStudy (gigi1103):

halp!!!

OpenStudy (gigi1103):

asap plz

OpenStudy (gigi1103):

@AlexandervonHumboldt2

OpenStudy (gigi1103):

@KendrickLamar2014

OpenStudy (gigi1103):

@Kitten_is_back

OpenStudy (gigi1103):

@pooja195

OpenStudy (gigi1103):

help plz!

OpenStudy (handsomechickenlover):

im trying to figure it out

OpenStudy (anonymous):

HELP ME PLEASE QUESTION ABOVE

OpenStudy (campbell_st):

a loss can be represented by what maths symbol... -, + , x or /

OpenStudy (anonymous):

9. (03.05 LC) The functions f(x) and g(x) are described below: f(x) = 32x + 8 g(x) = 32x −9 The graph of g(x) is obtained by shifting down the graph of f(x) by units. (5 points) Score: 0 of 5

OpenStudy (gigi1103):

ok so like x=-9+3+6

OpenStudy (handsomechickenlover):

-9+3+6?

OpenStudy (campbell_st):

that's correct

OpenStudy (campbell_st):

but you need an expression... so delete the x and equals symbol

OpenStudy (campbell_st):

Part B is asking you to get an answer for -9 + 3 + 6 = ?

OpenStudy (gigi1103):

is this for part b? or a

OpenStudy (gigi1103):

oh!!

OpenStudy (gigi1103):

so i just answer part b and thats it?

OpenStudy (gigi1103):

thanks sooo much!!

OpenStudy (campbell_st):

your statement x = -9 + 3 + 6 is the 1st part, but delete the x = just -9 + 3 + 6 Part B... get the value of -9 + 3 + 6

OpenStudy (gigi1103):

ok thank you you are amazing!!